40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 25 15:32:16 2019, gautam, Update, IOO, PMC and IMC locked again, some MEDM maintenance PMCctrl.png
    Reply  Wed Jun 26 10:55:13 2019, aaron, Update, IOO, PMC and IMC locked again, some MEDM maintenance 
       Reply  Wed Jun 26 11:11:40 2019, Milind, Update, IOO, PMC and IMC locked again, some MEDM maintenance burtgooey.pdf
    Reply  Wed Jun 26 18:28:24 2019, rana, Update, IOO, PMC and IMC locked again, some MEDM maintenance 
       Reply  Sun Jun 30 22:21:07 2019, Milind, Update, IOO, PMC and IMC locked again, some MEDM maintenance 
          Reply  Sun Jun 30 23:52:09 2019, Koji, Update, IOO, PMC and IMC locked again, some MEDM maintenance 
             Reply  Mon Jul 1 11:02:05 2019, Milind, Update, IOO, PMC and IMC locked again, some MEDM maintenance 
          Reply  Tue Jul 9 10:37:42 2019, Milind, Update, IOO, keyed psl crate, unstick.py, pmc autolocker code- working  
             Reply  Mon Jul 15 14:53:40 2019, Milind, Update, IOO, keyed psl crate, unstick.py 
                Reply  Mon Jul 15 18:55:05 2019, gautam, Update, IOO, Megatron hard-rebooted 
             Reply  Wed Jul 24 12:24:43 2019, Milind, Update, IOO, unstick.py and ifotest 
Message ID: 14699     Entry time: Wed Jun 26 10:55:13 2019     In reply to: 14696     Reply to this: 14700
Author: aaron 
Type: Update 
Category: IOO 
Subject: PMC and IMC locked again, some MEDM maintenance 

The PMC was locking again after Gautam's steps above. However, after I added the directional coupler between the mixer I and the servo card (coupled to the Agilent analyzer), the PMC was again not locking, except occasionally with gain of -10 dB.

I removed the coupler (so the mixer I goes directly to the PMC servo card, as Gautam had it), and the PMC was still not locking. While checking connections, I noticed that one of the SMA cables between the LO and the mixer was not even finger tight, so I tightened them to approximately the right torque with a non-torque wrench.

This did not lead to the PMC locking, so Millind helped me key the c1psl VME crate. I burt restored the latest snapshot. Now, the PMC locks up until gain of -5. I try burt restoring the previous snapshot, which was from when the PMC was locking, and now it locks. Adding in the directional coupler again leads to the PMC not locking, though this time removing the coupler restores the normal behavior. I also tried using the coupler with the coupling port connected to a 50 Ohm terminator, and this configuration also did not lock.

I had been using a ZFDC-20-5-S+ (0.1-2000 MHz) with SMA ports and SMA-to-BNC on the input and output ports (since the mixer has BNC connectors). To reduce the number of potentially flaky connections, I am trying the ZFDC-20-4 (1-1000 MHz) that I found with BNC ports. The PMC still doesn't lock.

To get some spectrum, I've connected the PMC servo card's 'mixer out' to the Agilent's A channel, and collected a spectra from [10 Hz, 75 kHz], [75 kHz, 750 kHz], and [750 kHz, 2 MHz].


Wed Jun 26 15:23:37 2019

After the lab cleaning, I added a BNC T on the mixer I port, so now the configuration is:

Mixer I -> BNC T

-> PMC Servo card FP1TEST

-> directional coupler -> coupled to the spectrum analyzer, out port is terminated with 50 Ohms.

I thought maybe the issue was that the TF from in->out on the directional coupler is not what I expect (and Gautam suggested the in-out port might block DC), but the PMC still does not lock in the above configuration, in which the coupler is not between the mixer and the servo board--so only reflections from the coupler should matter, I think.

However, even when I plug the mixer directly into the servo board, the PMC is not locking (again) with gain above -8 dB or so. I did a burt restore again, and this fixed the problem. I wasn't sure why this burt restore is working, because all I am changing is the DC output adjust voltage and the gain, and switching on/off FP1TEST. However, I observed that after running the PMC autolocking script, observing that the autolocker did not achieve lock as it swept through resonance, and cancelling the autolocker, the PMC again cannot be locked for high gains. When I let the autolocker complete, this doesn't happen, so probably I'm just not letting some channel return to its nominal value after being changed by the autolocker.

Now after another burt restore, I'm avoiding using the autolocker and am still having trouble locking with the BNC T + directional coupler configuration above. However, now I'm noticing that the PZT control mon is always railed, as long as FP1TEST is in the loop (and independent of the output adjust voltage). I try returning to the 'baseline' configuration (mixer -> PMC servo card directly), and the PMC locks but with only 0.68 V transmission (was >0.7 V before).


Per Gautam's earlier suggestion, I switched to using the Agilent 41800A probe instead of the directional coupler. I was able to lock the OMC with this probe on a BNC T coming out of the mixer (transmission is 0.71 V). I recorded the spectra of the PMC servo board's "Mixer Out" channel, and the mixer's I as seen by the probe. I recorded spectra from 10 Hz to 100 MHz. The soft linked netgpibdata folder I had in my users directory is no longer soft linked--presumably intentional so I don't tamper with it?

I'm a bit skeptical that I've used the probe correctly, so I'm checking out the manual.

Indeed, I needed to pull back the sheath; I also noticed that the GPIB script I've been using doesn't save the data from both channels when I take a spectrum in dual mode, so I'm taking the spectra again one at a time (lights are on, IMC is locked).

ELOG V3.1.3-